# Frustratingly Easy Test-Time Adaptation of Vision-Language Models

### v1

@@ 6 Limitations
Zero can seamlessly adapt a wide range of VLMs on arbitrary datasets without requiring extensive computational resources and is backed by theoretical justifications. However, we delineate three major limitations to our method which we report here. The first limitation concerns the preliminary observations which led to Zero, such as augmentation-induced overconfidence. For example, these observations may not persist if the VLM changes significantly in the future, potentially leading to poor adaptation. The second limitation stems from theoretical assumptions, the core one being the invariance of the marginal probability distribution to marginal entropy minimization. While our proposition guarantees invariance if entropy is minimized and the negative delta to the probability of the prevailing class is inferior to its default value, these theoretical assumptions may not hold all the times. In this work, we supported our assumptions with empirical verification but, as per the first limitation, these may not extend to the space of all models and datasets (as we have shown, for example, with satellite imagery). A third worthy-of-note limitation relates to the independence assumption among the views from which the marginal probability distribution is obtained. As we discussed in Section 2.3, the views themselves do not have any direct dependency, but they are still partially related through the source image from which they stem. Related to this, we particularly believe that extending Zero in a Retrieval-Augmented TTA setup can largely mitigate this limitation. Finally, despite being much lighter than the current state-of-the-art TTA strategies, Zero’s computational requirements in the visual branch scale linearly with the number of views, since all of them need to be independently forwarded. On this, we believe that exploring how to augment directly in the latent visual space to also circumvent the forward pass of the vision encoder is an intriguing direction.
@@ 7 Conclusions
We theoretically investigated Marginal Entropy Minimization, the core paradigm of the current research in Test-Time Adaptation with VLMs. Building on our theoretical insights, we introduced Zero: a frustratingly simple yet strong baseline for TTA, which only relies on a single batched forward pass of the vision encoder. Our experimental results on Natural Distribution Shifts and Fine-grained Classification unveil that Zero favorably compares to state-of-the-art TTA methods while requiring relatively negligible computation. We hope our findings will inspire researchers to push the boundaries of TTA further.
@@ A.2 Experimental verification
Appendix B Calibration and Overconfidence of CLIP on augmented Natural Distribution Shifts
@@
Figure 2: Expected Calibration Error (ECE) [7] of CLIP-ViT-B-16 across 5 datasets for robustness to natural distribution shifts. Blue is the ECE of zero-shot CLIP, orange is the ECE of zero-shot CLIP on an augmented version of the dataset after confidence-based thresholding.
@@
First, let us define the calibration of DNNs. Calibrating DNNs is crucial for developing reliable and robust AI systems, especially in safety-critical applications. A DNN is perfectly calibrated if the probability that its prediction is correct ($\hat{y}=y$) given a confidence score random variable $S$ is equal to its confidence score, which in our case is $s=\max\overline{p}(\cdot)$:
@@
These results are entirely consistent with Sec.3.1: the calibration error increases exclusively due to overconfidence, no matter the dataset. In parallel, the error rate of CLIP-ViT-B-16 can either remain close to its default value (e.g., ImageNet-Sketch), slightly decrease (e.g., ImageNet-R and -v2) or largely decrease (ImageNet-A).
Appendix C On the expected risk of $\overline{p}$ and $p$.
@@
Appendix D Tie breaking with Zero
A caveat of Zero are ties, i.e., cases with multiple classes having identical probability within the marginal probability distribution. This is clear to see when viewing Zero as its equivalent paradigm of voting among confident views, simply because more than one class may have an equal amount of “votes”. Throughout all the experiments of this work, ties are broken greedily. If a tie results from the top-10% views, the procedure for breaking it follows these two steps: ① sort the remaining 90% views by ascending entropy (most to least confident) and ② scan the views until a prediction is encountered that breaks the tie. Other than this, many alternative are possible, such as relying on the most confident prediction. Specifically, we have explored the following alternatives:
@@ Appendix D Tie breaking with Zero
Appendix E A failure mode for TTA: satellite imagery
@@
Table 5: TTA failures on EuroSAT.
[…]
CLIP-ViT-B-16 Zero-Shot 42.01
Ensemble 50.42
TPT 42.86
Zero 39.60
Zero+Ensemble 43.77
MaPLE Zero-Shot 48.06
TPT 47.80
PromptAlign 47.86
Zero 41.05
CLIP-ViT-B-16 + CLIP-ViT-L-14 Zero-Shot 54.38
[…]
Zero 42.74
@@
Nevertheless, we highlight an additional caveat about satellite imagery which is particularly detrimental for Zero, and relates to the base model error over augmentations. Recall that, in Zero, the usage of $\overline{p}$ is backed by theoretical motivations and the manual adaptation of the temperature is supported by two concurrent observations: augmentations-induced overconfidence and a comparable error rate between source and augmented images. Simply put, the latter condition is not verified for satellite imagery. To show this phenomenon, we follow the experimental setup of Sec.3.1 and examine the reliability diagrams of EuroSAT[10] and of its augmented counterpart in Figure 4. As per Section 3.1, we display the ECE and the Top1-Accuracy on each version of the dataset. From this perspective, one can note that the base model error largely increases, in this domain, when augmented views are present. The accuracy on source images is $42.01\%$, dropping to $35.21\%$ simply due to augmentations.
@@
Appendix F Additional Implementation Details

### v2

@@ 6 Limitations
Zero can seamlessly adapt a wide range of VLMs on arbitrary datasets without requiring extensive computational resources and is backed by theoretical justifications. However, we delineate four major limitations to our method which we report here.
Preliminary observations. The first limitation concerns the preliminary observations which led to Zero, such as augmentation-induced overconfidence or a comparable error rate between source and augmented datasets. These observations may not persist if VLMs or benchmarks change significantly in the future, potentially leading to poor adaptation. For example, we have observed a consistent failure case for TTA with EuroSAT [11], with Zero incurring large performance drops w.r.t. simple zero-shot classification. In Appendix F we unravel this worst-case further.
Theoretical assumptions. The second limitation stems from theoretical assumptions, the core one being the invariance of the marginal probability distribution to marginal entropy minimization. While our proposition guarantees invariance if entropy is globally minimized and the negative variation to the probability of the most probable class is less than the initial probability itself, these theoretical assumptions may not hold all the time. In this work, we supported our assumptions with empirical verification but, as per the first limitation, these may not extend to the space of all models and datasets. We refer the interested readers to Appendix A for a more in-depth discussion about the invariance of the prediction of $\overline{p}$ to MEM.
Independence among views. A third worthy-of-note limitation relates to the independence assumption among the views from which the marginal probability distribution is obtained. As we discussed in Section 2.3, the views themselves do not have any direct dependency, but they are still partially related through the source image from which they stem. Related to this limitation, we hypothesize that extending Zero in a Retrieval-Augmented TTA setup (or a cache-based one) could improve the results. The discussion on this topic is extended in Appendix H.
Linear complexity with respect to augmented views. Finally, despite being much lighter than the current state-of-the-art TTA strategies, Zero’s computational requirements in the visual branch scale linearly with the number of views, since all of them need to be independently forwarded. On this, we believe that exploring how to augment directly in the latent visual space to also circumvent the forward pass of the vision encoder is an intriguing direction.
@@ 7 Conclusions
We theoretically investigated Marginal Entropy Minimization, the core paradigm of the current research on Test-Time Adaptation with VLMs. Building on our theoretical insights, we introduced Zero: a frustratingly simple yet strong baseline for TTA, which only relies on a single batched forward pass of the vision encoder. Zero works by setting the temperature of the Softmax operator to “zero” before marginalizing across confident views, which is equivalent, in terms of output, to the widely known paradigm of majority voting. Our experimental results on Natural Distribution Shifts and Fine-grained Classification unveil that Zero favorably compares to state-of-the-art TTA methods while requiring relatively negligible computation. We hope our findings will inspire researchers to push the boundaries of TTA further.
Acknowledgements.
The authors acknowledge the CINECA award under the ISCRA initiative for the availability of high-performance computing resources and support. Matteo Farina is supported by the PRIN project “LEGO-AI” (Prot.2020TA3K9N) and the PAT project “AI@TN". This work was supported by the projects EU Horizon ELIAS (No. 101120237), AI4TRUST (No.101070190), FAIR - Future AI Research (PE00000013), funded by NextGeneration EU, and carried out in the Vision and Learning joint laboratory of Fondazione Bruno Kessler and the University of Trento, Italy.
@@ A.2 Experimental verification
A.3 Can invariance be anticipated?
In the proof of Proposition 2.1, we express the post-MEM embeddings as a function of the pre-MEM embeddings through a Taylor expansion. For this relationship to hold, the variation needs to be small. If the initial entropy is high, the gradients from MEM (and, thus, the variation between pre- and post-MEM embeddings) can be larger than what a Taylor expansion can accurately approximate. In such cases, Prop. 2.1 cannot be guaranteed.
Figure 2: Entropy of the pre-TTA marginal probability distribution vs the invariance ratio.
We execute a simple experiment using the validation set of ImageNet-1k, whose recipe is described below, to visualize this relationship.
We compute pre- and post-MEM marginal probability distributions. We sort the pre-MEM distributions in order of descending entropy (most to least uncertain) and quantize them into 10 bins. Bins shall be interpreted as follows: the leftmost bin contains the top 10% of samples with the highest entropy; the second bin contains samples outside the top-10% percentile but within the top-20%, and so on; the rightmost bin contains the bottom 10% of samples with the lowest entropy. For each bin we compute the invariance ratio, measuring how often the $\operatorname*{arg\,max}$ of the pre-MEM $\overline{p}$ does not change after MEM. Finally, we display a histogram with this data in Figure 2.
A trend appears: as the entropy decreases (left to right), invariance holds more and more often. Hence, intuitively, the most likely cases where invariance to MEM does not hold are those of high uncertainty in the initial marginal probability distribution. However, this may still be rare: even within the top 10% of most uncertain samples, invariance holds more than 82% of the time (leftmost bin).
Appendix B Additional Experiments: LAION-2B Pretraining, Context Optimization and Hyperparameter Inheritance
Table 5: Results on Natural Distribution Shifts when adapting CLIP-ViT-B-16 pretrained on the 2B English subset of LAION-5B. Top-1 accuracy is reported, and bold text indicates the best performer.
Method ImageNet A V2 R Sketch Mean
CLIP-ViT-B-16 (LAION2B)
Zero-Shot 69.27 37.08 61.27 78.83 54.85 60.26
Ensemble 70.43 38.32 62.28 80.41 55.54 61.40
TPT 70.61 41.94 62.96 80.40 55.48 62.28
Zero 71.39 48.71 63.53 80.59 55.82 64.01
Zero+Ensemble 72.14 49.02 64.32 82.42 56.53 64.89
Table 6: Results on Natural Distribution Shifts when adapting OpenAI’s CLIP-ViT-B-16, with CoOp-learned prompts. Top-1 accuracy is reported, and bold text indicates the best performer.
Method ImageNet A V2 R Sketch Mean
CoOp
Zero-Shot 71.51 49.71 64.20 75.21 47.99 61.72
TPT 73.64 57.77 66.72 78.03 49.56 65.14
Zero 74.12 61.57 67.15 78.43 49.77 66.21
Table 7: Fine-grained Classification with CLIP-ViT-B-16 pretrained on the 2B English subset of LAION-5B. Top-1 accuracy is reported, and bold text indicates the best performer.
Method FLWR DTD PETS CARS UCF CAL FOOD SUN AIR ESAT Mean Median
CLIP-ViT-B-16 (LAION2B)
Zero-Shot 69.71 54.43 89.37 89.94 64.02 95.82 81.38 70.60 26.04 47.05 68.84 70.16
Ensemble 68.70 54.55 87.76 89.98 67.64 96.51 81.64 70.62 25.68 49.64 69.27 69.66
TPT 69.47 54.53 89.00 90.72 66.68 96.16 81.76 71.34 26.73 48.81 69.52 70.41
Zero 70.82 55.20 89.77 91.95 67.23 96.13 83.65 71.21 28.25 45.01 69.92 71.02
Zero+Ensemble 68.01 55.95 87.67 91.87 69.11 96.54 83.83 71.09 28.10 47.10 69.93 70.10
This Appendix deals with enriching the experiments of Section 4, which focused on models officially released by OpenAI [28]. Here we focus on the comparison with TPT [38] and extend the analysis to: ① CLIP-ViT-B-16 pretrained on the 2B English Subset of LAION-5B [34]; ② OpenAI’s CLIP, transferred after supervised Context Optimization (CoOp) [56].
Implementation Details. For the experiments with LAION Pretraining, we use the open_clip repository, i.e., the official code for [2]. The pretrained keyword for this model is laion2b_s34b_b88k. For CoOp we use the context vectors learned on ImageNet-1k officially released by [56]. The experimental setup is analogous to Section 4 in all details. We do not tune any hyperparameters for these different initializations, but inherit them from the experiments with OpenAI models.
B.1 LAION-2B Pretraining
Table 5 reports experiments on Natural Distribution Shifts, from which we observe no differences w.r.t. OpenAI models: Zero largely outperforms TPT, and peak difference is reached with ImageNet-A [13]. Results on Fine-grained Classification are given in Table 7. We observe that Zero improves the zero-shot baseline better with this pretraining, and overcomes TPT with an average margin of $+0.4\%$. In contrast, ensembling textual prompts appears less effective. We speculate this is because the 7 templates were explicitly tuned and selected for OpenAI models. The worst-case scenario is confirmed with satellite imagery [11]; please refer to Appendix F for a deeper investigation.
B.2 Context Optimization (CoOp)
For this comparison, we follow [38] and report CoOp on Natural Distribution Shifts only, presenting results in Table 6. We further observe patterns consistent with OpenAI models, with Zero providing large improvements over TPT. Also here, the best-case scenario persists with ImageNet-A.
B.3 Hyperparameter Inheritance
Table 8: Natural Distribution Shifts (percentile = 0.1). TTA methods are grouped according to the baseline model and top-1 accuracy is reported. Bold text is the best method within each group.
Method ImageNet A V2 R Sketch Average
CLIP-ViT-B-16
Zero-Shot 66.73 47.87 60.86 73.98 46.09 59.11
Ensemble 68.34 49.89 61.88 77.65 48.24 61.20
TPT 68.98 54.77 63.45 77.06 47.94 62.44
MaPLe
Zero-Shot - 50.90 64.07 76.98 49.15 60.28
TPT - 58.08 64.87 78.12 48.16 62.31
PromptAlign - 59.37 65.29 79.33 50.23 63.55
Zero - 64.65$\pm$0.24 66.63$\pm$0.32 79.75$\pm$0.41 50.73$\pm$0.62 65.44
CLIP-ViT-B-16 + CLIP-ViT-L-14
ZeroShot 73.44 68.82 67.80 85.40 57.84 70.66
RLCF ${}_{t=3}^{\bm{\mathbf{t}_{\text{ctx}}}}$ 73.23 65.45 69.77 83.35 54.74 69.31
RLCF ${}_{t=3}^{\Theta_{v}}$ 74.85 73.71 69.77 86.19 57.10 72.32
Table 9: Finegrained classification (percentile = 0.1). Formatting follows other tables.
Method FLWR DTD PETS CARS UCF CAL FOOD SUN AIR ESAT Mean Median
CLIP-ViT-B-16
Zero-Shot 67.44 44.27 88.25 65.48 65.13 93.35 83.65 62.59 23.67 42.01 63.58 65.31
Ensemble 67.07 45.09 88.28 66.16 67.51 93.91 84.04 66.26 23.22 50.42 65.20 66.66
TPT 68.75 47.04 87.23 66.68 68.16 93.93 84.67 65.39 23.13 42.86 64.78 67.42
Zero 67.07 45.80 86.74 67.54 67.64 93.51 84.36 64.49 24.40 39.60 64.11 67.31
Zero+Ensemble 66.82 45.86 87.20 68.48 68.57 94.14 84.58 66.90 24.42 43.77 65.07 67.69
MaPLe
Zero-Shot 72.23 46.49 90.49 65.57 68.69 93.53 86.20 67.01 24.74 48.06 66.30 67.85
TPT 72.37 45.87 90.72 66.50 69.19 93.59 86.64 67.54 24.70 47.80 66.49 68.37
PromptAlign 72.39 47.24 90.76 68.50 69.47 94.01 86.65 67.54 24.80 47.86 66.92 68.98
Zero 71.20 47.70 90.17 67.91 69.49 94.12 86.78 67.55 25.57 41.05 66.15 68.70
CLIP-ViT-B-16 + CLIP-ViT-L-14
ZeroShot 75.76 51.83 92.86 76.16 73.70 94.04 88.03 66.96 30.54 54.38 70.43 74.73
RLCF ${}_{t=1}^{\bm{\mathbf{t}_{\text{ctx}}}}$ 71.58 50.34 89.01 69.76 69.84 94.09 85.90 67.33 23.71 46.87 66.84 69.80
RLCF ${}_{t=3}^{\bm{\mathbf{t}_{\text{ctx}}}}$ 72.49 51.93 89.55 72.91 72.31 95.00 86.84 69.04 25.40 45.96 68.14 72.40
RLCF ${}_{t=1}^{\Theta_{v}}$ 72.56 52.21 89.51 63.12 71.49 94.65 86.90 68.50 24.06 47.74 67.07 70.00
RLCF ${}_{t=3}^{\Theta_{v}}$ 71.74 53.27 91.15 70.93 73.24 94.73 87.28 69.38 28.54 47.41 68.77 71.34
Zero 75.34 54.22 92.90 77.33 74.26 94.52 87.57 68.05 32.11 42.74 69.90 74.80
In all experiments so far, including Section 4 as well as Tables 5, 6 and 7, we employed a percentile for confidence-based filtering set to 0.3. This value was obtained after validation on ImageNet-1k with OpenAI’s CLIP-ViT-B-16 and kept fixed for all models and datasets. Here, we show that Zero obtains favorable performance even if the percentile for confidence-based filtering is not tuned in any way, but set to 0.1 by “inheriting” the value used in TPT [38]. These results are given in Tables 8 and 9. Surprisingly, some datasets within the Natural Distribution Shifts benchmark benefit from this more restrictive filtering (ImageNet-A above all), while we observe that Finegrained classification tends to improve when more views are retained. The core findings, however, are entirely unchanged: the best case remains ImageNet-A, the worst-case remains EuroSAT, and Zero outperforms competitor in most datasets, no matter the experimental setup.
@@ B.3 Hyperparameter Inheritance
Appendix C Calibration and Overconfidence of CLIP on augmented Natural Distribution Shifts
@@
Figure 3: Expected Calibration Error (ECE) [8] of CLIP-ViT-B-16 across 5 datasets for robustness to natural distribution shifts. Blue is the ECE of zero-shot CLIP, and orange is the ECE of zero-shot CLIP on an augmented version of the dataset after confidence-based thresholding.
@@
First, let us define the calibration of DNNs. Calibrating DNNs is crucial for developing reliable and robust AI systems, especially in safety-critical applications. A DNN is perfectly calibrated if the probability that its prediction is correct ($\hat{y}=y$) given a confidence score random variable $S$ is equal to its confidence score. The confidence score is commonly taken as the maximum of the output probability vector of the model, i.e., $s=\max p(\cdot)$:
@@
Figure 5: Reliability diagram (10 bins) for CLIP-ViT-B-16 pretrained on LAION-2B when transferred zero-shot on ImageNet-1k. (left) Source Dataset, (right) Augmented version of the dataset.
These results are entirely consistent with Sec.3.1: the calibration error increases exclusively due to overconfidence, no matter the dataset. In parallel, the error rate of CLIP-ViT-B-16 can either remain close to its default value (e.g., ImageNet-Sketch), slightly decrease (e.g., ImageNet-R and -v2) or largely decrease (ImageNet-A). We observe an identical pattern for CLIP models pretrained on LAION. For reference, see Figure 5.
Appendix D On the expected risk of $\overline{p}$ and $p$.
@@
Appendix E Tie breaking with Zero
A caveat of Zero are ties, i.e., cases with multiple classes having identical probability within the marginal probability distribution. This is clear to see when viewing Zero as its equivalent paradigm of voting among confident views, simply because more than one class may have an equal amount of “votes”. Throughout all the experiments of this work, ties are broken greedily. If a tie results from the top views, the procedure for breaking it follows these two steps: ① sort the remaining views by ascending entropy (most to least confident) and ② scan the views until a prediction is encountered that breaks the tie. Other than this, many alternative are possible, such as relying on the most confident prediction. Specifically, we have explored the following alternatives:
@@ Appendix E Tie breaking with Zero
Appendix F A failure mode for TTA: satellite imagery
In our experiments, we find that an extremely OOD domain represents a consistent failure mode for TTA: satellite imagery. In all comparison groups, a zero-shot baseline largely outperforms any TTA strategy when evaluated on EuroSAT[11]:
  • in Group 1 the zero-shot baseline CLIP-Ensemble largely outperforms the best TTA strategy Zero+Ensemble;
  • in Group 2, zero-shot MaPLe outperforms PromptAlign;
@@ Appendix F A failure mode for TTA: satellite imagery
Here, we qualitatively and quantitatively report two main root causes for failures.
@@
Nevertheless, we highlight an additional caveat about satellite imagery which is particularly detrimental for Zero, and relates to the base model error over augmentations. Recall that, in Zero, the usage of $\overline{p}$ is backed by theoretical motivations, and the manual adaptation of the temperature is supported by two concurrent observations: augmentations-induced overconfidence and a comparable error rate between source and augmented images. Simply put, the latter condition is not verified for satellite imagery. To show this phenomenon, we follow the experimental setup of Sec.3.1 and examine the reliability diagrams of EuroSAT[11] and of its augmented counterpart in Figure 6. As per Section 3.1, we display the ECE and the Top1-Accuracy on each version of the dataset. From this perspective, one can note that the base model error largely increases, in this domain, when augmented views are present. The accuracy on source images is $42.01\%$, dropping to $35.21\%$ simply due to augmentations.
@@
Appendix G Natural Distribution Shifts vs Fine-grained Classification
Throughout the manuscript, one can observe that Zero consistently provides larger improvements in Natural Distribution Shifts than it does in the Finegrained suite. We thus devote this section to digging deeper into this matter.
Perhaps unsurprisingly, we posit that Zero improves over the zero-shot baseline if the zero-shot error rate of the model does not largely increase with augmented views. As Fig.1(b) displays, this is the case for all Natural Distribution Shifts datasets. To understand any different behaviors, we repeat the same experiment of Section 3.1 for the entire Fine-grained suite and report the results in Table 10.
Table 10: Comparison among ① CLIP’s zero-shot accuracy, ② CLIP’s accuracy on the augmented counterpart of the dataset, and ③ Zero. The augmented datasets are crafted following the protocol of Section 3.1 . “Gap” is defined as CLIP’s zero shot accuracy minus its accuracy on the augmented dataset. “Improvement” is defined as the accuracy of Zero minus that of zero-shot CLIP. Spearman’s coefficient between “Gap” and “Improvement” equals $-0.95$: as the “Gap” decreases (i.e., the lower the error on augmented views) Zero provides more substantial improvements.
Method FLWR DTD PETS CARS UCF CAL FOOD SUN AIR
① Zero-Shot 67.44 44.27 88.25 65.48 65.13 93.35 83.65 62.59 23.67
② Augmented 66.19 44.90 86.17 65.88 65.59 92.62 83.25 62.97 23.52
Zero (perc = $0.1$) 67.07 45.80 86.74 67.54 67.64 93.51 84.36 64.49 24.40
Gap $=$-- $+$1.25 --0.63 $+$2.08 --0.40 --0.46 $+$0.73 $+$0.40 --0.38 $+$0.15
Improvement $=$-- --0.37 $+$1.53 --1.51 $+$2.06 $+$2.51 $+$0.16 $+$0.71 $+$1.90 $+$0.73
Please note that, in the table, the percentile for confidence-based in filtering in Zero is set to $0.1$, since the protocol for generating the augmented datasets follows the setup of TPT, which also uses a cutoff percentile of $0.1$, and that we omit EuroSAT since an analogous experiment was presented in the previous Appendix.
Overall, we observe a strong correlation between the error gap and the improvement provided by Zero, with Spearman’s coefficient being $-0.95$ across datasets. This result shows that the correlation is negative, i.e., the lower the error gap, the larger the improvement (or, in other words, the better the zero-shot performance on augmented views, the larger the improvement of Zero). This pattern is also consistent with the experiments on EuroSAT reported in the previous Appendix. Understanding why augmentations induce larger or smaller errors may be a case-by-case matter that relates to the nature of the datasets. Here, we pinpoint two possible reasons:
  • The semantic space of the ImageNet variants of the Natural Distribution Shifts benchmark comprises many common categories, which may have appeared frequently during CLIP’s pretraining. Hence, it seems reasonable that CLIP is robust w.r.t. augmented views of images belonging to these categories. In the Fine-grained classification suite, datasets such as SUN397 and Caltech101 also contain common object categories, which is consistent with the results shown above. In contrast, other datasets such as Flowers102 and Oxford-Pets span much less frequent concepts.
  • Other than the semantic classification space, images’ visual appearance also plays an important role. For example, datasets such as FGVC-Aircraft and Stanford Cars still contain rare concepts, but Zero largely improves over the baseline nonetheless. Our augmentation setup is simple, and only contains random resized crops and random horizontal flips, which can constitute a “zoom-in” to a random portion of the image. For some benchmarks, this is useful as it may trigger CLIP’s capabilities to recognize small details, such as logos, or even reading text, such as the car brand or the airline name. In contrast, more object-centric datasets such as Flower102, may lead to missing precious visual features (e.g., the stem).
In our work we did not search for the best data augmentations but rather stuck to an established setting, using the same augmentations setup for all datasets. Nevertheless, the performance of Zero is linked to the impact that data augmentations have on how the model perceives images, and we believe this is an interesting research direction to pursue.
Appendix H Independence among views in the setup of Test-Time Adaptation
The theoretical framework of Section 2.3 models an ideal scenario, where independence holds among different inputs. To clarify, this means that the model’s error on view $\mathbf{x}_{i}$ should not be correlated with the error on any other view $\mathbf{x}_{j}$, which allows writing the compound error with a binomial distribution as in (6).
In practice, achieving perfect independence is challenging, if not impossible. Hence, a suitable approximation strategy to mitigate this issue is to promote diversity. In classical ensembling theory, a well-established approach is to train different models on different subsets of the available data. Similarly, the augmentation scheme of random cropping aligns with this approach by presenting the model with different portions of the image each time.
Moreover, ideally, the augmentation pipeline should not change the underlying label of the original input and guarantee that the model’s error rate on augmented views remains comparable to the error rate on the original inputs belonging to the same category. In practice, this entails that augmentations should not disrupt the visual appearance of the image, and, consequently, some views may result in a slight or moderate correlation, because some “parts” of the source image will overlap among them. An analogy with classical literature can be drawn also in this case. Specifically, when not enough data are available, overlaps among the training sets of different models are required to ensure convergence. Consequently, models producing slightly or moderately correlated predictions are more likely to emerge.
Appendix I Additional Implementation Details